# Do Large Language Models Have an English “Accent”? Evaluating and Improving the Naturalness of Multilingual LLMs

### v2

@@ 2 Related Work
Translationese detection, a well-established task in machine translation, aims to determine whether a text is originally written in the target language or translated from another language (Volansky et al., 2015; Wintner, 2016). For instance, Freitag et al. (2022) employ a range of linguistic features – including type-token ratio, lexical density, answer lengths, dependency tree height, constituency tree height, perplexity, etc. – to train a classifier to distinguish between machine-translated and naturally occurring sentences. However, these classifiers are prone to overfit on specific training data. In an effort to enhance the naturalness of machine translation outputs, Freitag et al. (2022) also tagged parallel training data based on target-side naturalness, contrasting models trained on natural versus translated text. Our approach builds on the concept of contrasting natural versus unnatural texts but avoids reliance on pre-trained classifiers. Instead, we use automatically manipulated unnatural texts and preference learning, extending the analysis beyond machine translation to general multilingual text generation.
Linguistic diversity evaluation (Tevet and Berant, 2021) is another area closely related to naturalness, as a key marker of unnaturalness in synthetic text is reduced linguistic diversity (Vanmassenhove et al., 2021; Guo et al., 2024). Our work draws inspiration from the way these diversity features are computed. Past work compares the diversity of generated texts to human texts and consider texts to be natural if they approach the human level of diversity (Freitag et al., 2022). However, such an approach primarily assesses the spread of the distribution while overlooking more holistic comparisons of linguistic features. To address this, we introduce metrics that directly compare vocabulary and syntactic distributions between human and machine-generated texts.
@@ 3.2 Syntactic Naturalness
To compute the structural similarity between pairs of dependency trees, we propose to use the Weisfeiler-Lehman (WL) graph kernel (Shervashidze et al., 2011). The WL kernel iteratively refines node labels based on the labels of neighboring nodes, thereby constructing a hierarchical encoding of the graph structure. More formally, let $T_{1}$ and $T_{2}$ be two dependency trees with respective vertex, i.e., word, sets $V_{1}$ and $V_{2}$. The WL kernel, $K_{\text{WL}}(T_{1},T_{2})$, measures the similarity between $T_{1}$ and $T_{2}$ by comparing their subtree structures. Initially, each node in the trees is assigned a label based on its original POS tag. Then, at each iteration $h$, the label of each node is updated based on the multiset of labels of its neighboring nodes, and the new labels are hashed into unique labels. This process continues for $H$ iterations, and the kernel value is computed as the number of matching node labels across all iterations:
@@
To conclude, our proposed syntactic naturalness metric quantifies the syntactic divergence between human and model-generated sentences by examining the distribution of their dependency tree structures. This approach considers both the dependency relationships and the hierarchical arrangement of substructures at multiple levels.
@@ 4.1 Dataset and Evaluation
Table 2: Extracted syntactic patterns that show the influence of English structures on how Llama-3.1 generates Chinese text. All frequencies are based on 40,000 $n$-grams of the same $n$. The Chinese and English outputs provided are not semantically aligned nor translations.
@@ 4.2 Multilingual LLMs
We experiment with three families of open-weight LLMs: Llama (Dubey et al., 2024), Qwen (Yang et al., 2024), and Mistral (Jiang et al., 2023). These models were selected for their state-of-the-art performance on diverse English and multilingual benchmarks. Additionally, they are developed by teams from regions where English, Chinese, and French are the official languages, respectively. While the exact composition of the training data for each model is not publicly available, we speculate that all are predominantly English-centric, though we expect Qwen to include more Chinese data and Mistral to include more French data than the others. We benchmark the two most recent versions from each model family to track changes in naturalness performance over time. For all models, we test their moderately scaled versions, using open-source implementations from the Transformers library (Wolf et al., 2020). We focus on the instructed versions of the models, as these are more commonly used in real-world applications compared to the base versions.
@@ 4.3 Results and Analysis
The results on our naturalness dataset according to our new metrics (Section 3) are presented in Table 1. Unsurprisingly, the results demonstrate a consistent improvement in naturalness across newer versions of LLMs compared to earlier models (e.g., from Qwen-1.5 to Qwen-2). However, our metrics highlight a noticeable gap between the level of naturalness of human-written and LLM-generated text, particularly for non-English generations, validating our initial hypothesis. Note that the human reference value is calculated by randomly selecting non-overlapping subsets of human-written texts and computing the divergence between them.
@@
In terms of lexical divergence, English outputs – especially from Mistral-v0.3 – approach human reference values, while larger gaps persist in Chinese and French, with Chinese showing the most significant difference. This indicates that the models are lexically less natural in languages other than English, and the naturalness gap is more pronounced for languages that are typologically distant from English.
@@
When examining syntactic divergence, all models in all languages show significant differences from the human reference values. Since dependency trees, when parsed using Universal Dependencies grammar, share a common space across languages, we can perform cross-lingual comparisons of syntactic structures. In Figure 1, we show Llama-3.1’s English “accent” in syntax when generating Chinese outputs. Despite the fact that Llama-3.1 is closer to human reference than the other LLMs (as shown in Table 1), it is easy to observe how Llama-3.1’s syntactic structures are more aligned with human-written English than human-written Chinese descriptions are, suggesting that English linguistic patterns influence Llama-3.1’s Chinese generations. We can observe a similar effect in French, even if less noticeable since English and French belong to linguistic families (West Germanic and Romance, respectively) that are closer. The English accent of Llama-3.1 can also be seen in our case study of unnatural language patterns in text generated by Llama-3.1 in Chinese. Table 2 presents examples of POS tag n-grams that frequently appear in both English and Llama-3.1-generated Chinese, but much less frequently in native Chinese texts.
@@
Impact of prompting language. For the previous benchmark results, we prompt the models in the same language in which we expect them to generate. To evaluate whether the input language affects the naturalness of the generations, we conduct an ablation study in a cross-lingual setting. Specifically, we manually translate the Chinese and French prompts into English and feed these English prompts into the LLM, while still instructing the LLM to generate responses in Chinese or French, that is, only the instructions are altered. The results for the Chinese and French generations are presented in Figures 2 and 3 in the Appendix. For Chinese outputs, prompting in Chinese consistently leads to better naturalness. However, for French outputs, the results vary across models. We hypothesize that this is influenced by the proportion of instruction-response pairs the models encountered in different languages during instruction tuning. Additionally, the closer typological relationship between French and English, compared to Chinese, may also contribute.
@@
Essay Generation Lexical Divergence $\downarrow$ 34.89 35.06 34.56 41.45 40.58 40.60
@@
Question Answering Lexical Divergence $\downarrow$ 28.05 28.01 28.04 28.09 27.58 26.92
@@
夹乒乓大赛 The Ping-Pong Clamping Contest Unaligned
@@ 5.1 Preference Dataset Construction
The original instructions from the SFT datasets are preserved, with the initial response serving as the preferred response. To generate the unnatural, i.e., rejected, response we apply synthetic manipulations through paraphrasing and back-translation via English. The prompts we use to generate the rejected responses are listed in Appendix D. We ensure semantic consistency by using BLEU (Papineni et al., 2002) to filter out pairs of chosen and rejected responses with insufficient similarity. Moreover, to ensure that the linguistic style of the rejected responses differs significantly from the chosen ones, we also filter out pairs with overly high BLEU scores. Based on an empirical analysis of BLEU score distributions, we define the final threshold as $0.15<\textsc{Bleu}(\text{Chosen},\text{Rejected})<0.9$. We also filter out responses shorter than 10 words (often in the style of multiple-choice questions), as they do not convey enough lexical or syntactic style. The statistics of our two resulting preference datasets are presented in Table 7.
@@ 5.2 Experimental Setup
We believe that it is beneficial to perform naturalness alignment on models that have already undergone SFT and previous rounds of preference tuning, since preference tuning is typically done iteratively (Dubey et al., 2024). We conduct experiments with the instruction-tuned versions of Llama-3.1 and Qwen2, which demonstrated the highest and lowest performance, respectively, in naturalness for Chinese in Table 1. To increase GPU memory usage efficency and to avoid catastrophic forgetting, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2022) in conjunction with DPO. Detailed hyperparameters for both LoRA and DPO are provided in Appendix E.
We explore two approaches: (1) generating the preference-tuning dataset with the same model that will be aligned (e.g., using Llama-3.1 to create training data for Llama-3.1), and (2) cross-model alignment, where the dataset generated by one model is used to align another model (e.g., using Llama-3.1 to create training data for Qwen2).
@@ 5.3 Results
Although we experiment with generating rejected responses with both paraphrasing and back-translation, the latter consistently outperforms the former by a small margin. We believe this is due to back-translation being more effective at introducing translationese artifacts into the rejected responses. Therefore, the final results in Table 3 are based on back-translation. However, there is no definitive conclusion on whether rejected responses should be generated using the same model undergoing alignment or a different model, and future work could explore generating response pairs with a combination of different models.
In addition, we evaluate our aligned models on the Chinese Massive Multitask Language Understanding (CMMLU) benchmark (Li et al., 2023) to ensure that their general capabilities are not compromised after naturalness alignment. Our results show that our alignment method not only improves linguistic naturalness but also slightly enhances overall language understanding performance. Interestingly, we observe that naturalness does not always positively correlate with language understanding performance. For example, Qwen2, despite achieving much higher scores than Llama-3.1 on the CMMLU benchmark, demonstrates lower naturalness, probably due to the heavy use of synthetic data. This highlights the need for future research to consider naturalness as a complementary metric alongside conventional benchmark scores.
@@ 6 Conclusion
In this paper, we address the naturalness challenges of language models, particularly in multilingual contexts, by conducting experiments in both French and Chinese. We introduce two corpus-level metrics to quantify the naturalness of model output distributions: one focused on vocabulary (lexical naturalness) and the other on grammatical structure (syntactic naturalness). These metrics are interpretable and free from biases introduced by external embeddings. Using these metrics, we benchmark state-of-the-art multilingual LLMs and analyze how factors such as training data, prompting language, and decoding strategy influence the generated language. Our analysis provides valuable insights into the current multilingual LLM landscape, complementing traditional performance benchmarks for task-solving capabilities. Finally, we propose an alignment method using DPO and a synthetically manipulated preference dataset to enhance the naturalness of model outputs. Experiments show that our aligned models consistently improve in both lexical and syntactic naturalness.
@@
Our experiments do not include any low-resource languages because our approach relies on the availability of ground-truth distributions from native, human-written data, which is often scarce or unavailable for many languages. While our naturalness evaluation metrics are designed to be language-agnostic, they depend on reliable word tokenizers and dependency parsers for the languages studied. Unfortunately, such tools are still lacking for most low-resource languages. However, we argue that naturalness evaluation should not be the priority for these languages at this stage, as it is only meaningful once models achieve a certain baseline level of overall performance.
@@ B.1 Prompts
Target
Language
Prompt
Language
Prompt
English English Please write a summary description of the follwing Wikipedia entry in English: <wiki entry in English>
Chinese Chinese {CJK}UTF8gbsn请用简体中文写出以下维基百科词条的摘要描述:<wiki entry in Chinese>
English Please write a summary description of the follwing Wikipedia page in Simplified Chinese: <wiki entry in Chinese>
French French Veuillez rédiger une description résumée de l’entrée Wikipedia suivante en français: <wiki entry in French>
English Please write a summary description of the following Wikipedia entry in French: <wiki entry in French>
Table 5: Prompts for Wikipedia description generation. For generations in Chinese and French, we experiment with both prompts in the target language and in English, to study the impact of prompting language on naturalness.
@@ Appendix C Implementation of Naturalness Metrics
Appendix D Preference Dataset Generation
@@
D.1 Prompts
Transformation Prompt
Back Translation zh->en
You are a professional translator. You always show the translated version, without any additional explanations or format changes.
Translate from English into Simplified Chinese:
<text in Chinese>
en->zh
You are a professional translator. You always show the translated version, without any additional explanations or format changes.
Translate from Simplified Chinese into English:
<text in English>
Paraphrasing
You are a professional editor who revises word choices and restructures sentences while preserving the original meaning. You
always show the edited version, without any additional explanations or format changes.
Edit the following text:
<text in Chinese>
Table 6: Prompts used to generate rejected responses in the preference tuning dataset.
@@ D.1 Prompts
D.2 Dataset Statistics
#Train #Test Lenprompt Lenchosen Lenrejected BLEU($chosen,rejected$)
Essay 10,050 4,950 29 564 523 0.29
QA 4,281 2,109 52 185 176 0.46
Table 7: Statistics of our preference tuning datasets constructed with back translation. All tokenization for length calculation was performed using the Llama-3.1 tokenizer.
@@ D.2 Dataset Statistics
Appendix E Hyperparameters for Preference Tuning
General LORA DPO
Parameter learning_rate max_grad_norm warmup_ratio per_device_batch_size lora_alpha lora_dropout r target_modules beta
Value 5e-6 0.3 0.1 6 128 0.05 256 all-linear 0.5
Table 8: Hyperparameters used for preference tuning. Initial values are based on recommendations from the DPO and LoRA papers, with minimal additional tuning applied.
@@ Appendix E Hyperparameters for Preference Tuning
Appendix F Additional Results
Additional results analyzing the factors influencing model naturalness are provided in Figures 2, 3, and 4.

### v3

@@ 2 Related Work
Translationese detection, a well-established task in machine translation, aims to determine whether a text is originally written in the target language or translated from another language (Volansky et al., 2015; Wintner, 2016). For instance, Freitag et al. (2022) employ a range of linguistic features – including type-token ratio, lexical density, answer lengths, dependency tree height, constituency tree height, perplexity, etc. – to train a classifier to distinguish between machine-translated and naturally occurring sentences. However, these classifiers are prone to overfit on specific training data. In an effort to enhance the naturalness of machine translation outputs, Freitag et al. (2022) also tagged parallel training data based on target-side naturalness, contrasting models trained on natural versus translated text. Our approach builds on the concept of contrasting natural versus unnatural texts but avoids reliance on pre-trained classifiers. Instead, we use automatically manipulated unnatural texts and preference learning. We also extend the analysis beyond machine translation to general multilingual text generation.
Linguistic diversity evaluation (Tevet and Berant, 2021) is another area closely related to naturalness, as a key marker of unnaturalness in synthetic text is reduced linguistic diversity (Vanmassenhove et al., 2021; Guo et al., 2024). Our work draws inspiration from the way these diversity features are computed. Past work compares the diversity of generated texts to human texts and consider texts to be natural if they approach the human level of diversity (Freitag et al., 2022). However, such an approach primarily assesses the dispersion of the distribution while overlooking more holistic comparisons of linguistic features. To address this, we introduce metrics that directly compare vocabulary and syntactic distributions between human and machine-generated texts.
@@ 3.2 Syntactic Naturalness
To compute the structural similarity between pairs of dependency trees, we propose to use the Weisfeiler-Lehman (WL) graph kernel (Shervashidze et al., 2011). The WL kernel iteratively refines node labels based on the labels of neighboring nodes, thereby constructing a hierarchical encoding of the graph structure. More formally, let $T_{1}$ and $T_{2}$ be two dependency trees with respective vertex, i.e., word, sets $V_{1}$ and $V_{2}$. The WL kernel, $K_{\text{WL}}(T_{1},T_{2})$, measures the similarity between $T_{1}$ and $T_{2}$ by comparing their subtree structures. Initially, each node in the trees is assigned a label based on its original POS tag. Then, at each iteration $h$, we aggregate the labels of the node’s neighbors into a multiset, which is then hashed into a new unique label. This process continues for $H$ iterations, and the kernel value is computed as the number of matching node labels across all iterations:
@@
To conclude, our proposed syntactic naturalness metric quantifies the syntactic divergence between human and model-generated sentences by examining the distribution of their dependency tree structures. This approach considers both the dependency relationships and the hierarchical arrangement of substructures at multiple levels. Each step in our approach (e.g., the POS tagger, WL kernel, and MMD) has been externally validated for accuracy and effectiveness.
@@ 4.1 Dataset and Evaluation
Table 2: Extracted syntactic patterns that show the influence of English structures on how Llama-3.1 generates Chinese text. All frequencies are based on 40,000 $n$-grams of the same $n$. The Chinese and English outputs provided are neither semantically aligned nor translations.
@@ 4.2 Multilingual LLMs
We experiment with three families of open-weight LLMs: Llama (Dubey et al., 2024), Qwen (Yang et al., 2024), and Mistral (Jiang et al., 2023). These models are selected for their state-of-the-art performance on diverse English and multilingual benchmarks. Additionally, they are developed by teams from regions where English, Chinese, and French are the official languages, respectively. While the exact composition of the training data for each model is not publicly available, we speculate that all are predominantly English-centric, though we expect Qwen to include more Chinese data and Mistral to include more French data than the others. We benchmark the two most recent versions from each model family to track changes in naturalness performance over time. For all models, we test their moderately scaled versions, using open-source implementations from the Transformers library (Wolf et al., 2020). We focus on the instruction-tuned versions of the models, as these are more commonly used in real-world applications compared to the base versions.
@@ 4.3 Results and Analysis
Table 1 presents the naturalness performance of all models based on our proposed metrics (introduced in Section 3). As expected, the results show a consistent improvement in naturalness across newer versions of LLMs compared to their earlier counterparts (e.g., from Qwen-1.5 to Qwen-2). However, our metrics also reveal a persistent gap between the naturalness of human-written and LLM-generated texts, especially in non-English outputs, supporting our initial hypothesis.
Note that the human reference value is computed by randomly selecting non-overlapping subsets of human-written texts and measuring the divergence between them. These values are not obtained under the exact same conditions as the human-model divergences, which may involve overlapping prompts. This stricter setup makes the human reference value we report an upper bound on the human divergence baseline. As we show below, even this upper bound remains substantially lower than the human-model divergences.
@@
In terms of lexical divergence, English outputs – especially from Mistral-v0.3 – approach human reference values, while larger gaps persist in Chinese and French, with Chinese showing the most significant difference. This indicates that the models are lexically less natural in languages other than English, and the naturalness gap is more pronounced for languages that are typologically more distant from English. The syntactic divergence values across different languages are not directly comparable. This is due to the fact that the UD grammar parses languages at varying levels of granularity, as evidenced by the lower human divergence value in Chinese compared to English or French.
@@
When examining syntactic divergence, all models in all languages show significant differences from the human reference values. Since dependency trees, when parsed using the UD grammar, share a common space across languages, we can perform cross-lingual comparisons of syntactic structures. In Figure 1, we show Llama-3.1’s English “accent” in syntax when generating Chinese outputs. Although Llama-3.1 is overall closer to the human distribution than other LLMs (as shown in Table 1), its syntactic structures still exhibit greater similarity to human-written English than to native Chinese, suggesting that English syntactic patterns influence its Chinese generations. We can observe a similar effect in French, even if less noticeable since English and French belong to linguistic families (West Germanic and Romance, respectively) that are closer. The English accent of Llama-3.1 can also be seen in our case study of unnatural language patterns in text generated by Llama-3.1 in Chinese. Table 2 presents examples of POS tag n-grams that frequently appear in both English and Llama-3.1-generated Chinese, but much less frequently in native Chinese.
@@
Impact of prompting language. For the previous benchmark results, we prompt the models in the same language in which we expect them to generate. To evaluate whether the input language affects the naturalness of the generations, we conduct an ablation study in a cross-lingual setting. Specifically, we manually translate the Chinese and French prompts into English and feed these English prompts to the LLM, while still instructing it to generate responses in Chinese or French. In other words, only the instructions are modified, the target language of the output remains unchanged. The results for the Chinese and French generations are presented in Figures 2 and 3. For Chinese outputs, prompting in Chinese consistently leads to better naturalness. However, for French outputs, the results vary across models. We hypothesize that this is influenced by the proportion of instruction-response pairs the models encountered in different languages during instruction tuning. Additionally, the closer typological relationship between French and English, compared to Chinese, may also contribute.
@@
Essay
Generation
Lexical Divergence $\downarrow$ 34.89 35.06 34.56 41.45 40.58 40.60
@@
Question
Answering
Lexical Divergence $\downarrow$ 28.05 28.01 28.04 28.09 27.58 26.92
@@
夹乒乓大赛
The Ping-Pong
Clamping Contest
Unaligned
@@ 5.1 Preference Dataset Construction
[…]
@@ 5.2 Experimental Setup
We believe that it is beneficial to perform naturalness alignment on models that have already undergone SFT and previous rounds of preference tuning, since preference tuning is typically done iteratively (Dubey et al., 2024). We conduct experiments with the instruction-tuned versions of Llama-3.1 and Qwen2, which demonstrated the highest and lowest performance, respectively, in naturalness for Chinese in Table 1. To increase GPU memory usage efficiency and to avoid catastrophic forgetting, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2022) in conjunction with DPO. Detailed hyperparameters for both LoRA and DPO are provided in Appendix F.
We explore two approaches: (1) self-alignment, where the preference-tuning dataset is generated by the same model being aligned (e.g., using Llama-3.1 to create training data for Llama-3.1), and (2) cross-model alignment, where one model generates the dataset used to align a different model (e.g., using Llama-3.1 to produce training data for Qwen2).
@@ 5.3 Results
Although we experiment with generating rejected responses using both paraphrasing and back-translation, the latter consistently yields slightly better performance than the former. We believe this is due to back-translation being more effective at introducing translationese artifacts into the rejected responses. Therefore, the final results in Table 3 are based on back-translation. However, there is no definitive conclusion on whether rejected responses should be generated using self-alignment or cross-model alignment, and future work could explore generating response pairs with a combination of different models.
In addition, we evaluate our aligned models on the Chinese Massive Multitask Language Understanding (CMMLU) benchmark (Li et al., 2023) to ensure that their general capabilities are not compromised after naturalness alignment. Our results show that our alignment method not only improves linguistic naturalness but also slightly enhances overall language understanding performance. Interestingly, we observe that naturalness does not always positively correlate with language understanding performance. For example, Qwen2, despite achieving much higher scores than Llama-3.1 on the CMMLU benchmark, demonstrates lower naturalness, potentially due to the heavy use of synthetic data. This highlights the need for future research to consider naturalness as a complementary metric alongside conventional benchmark scores.
@@ 6 Conclusion
In this paper, we address the naturalness challenges of LLMs, particularly in multilingual contexts, by conducting experiments in English, French and Chinese. We introduce two corpus-level metrics to quantify the naturalness of model output distributions: one focused on vocabulary (lexical naturalness) and the other on grammatical structure (syntactic naturalness). These metrics are interpretable and free from biases introduced by external embeddings. Using these metrics, we benchmark state-of-the-art multilingual LLMs and analyze how factors such as training data, prompting language, and decoding strategy influence the generated language. Our analysis provides valuable insights into the current multilingual LLM landscape, complementing traditional performance benchmarks for task-solving capabilities. Finally, we propose an alignment method using DPO and a synthetically manipulated preference dataset to enhance the naturalness of model outputs. Experiments show that our aligned models consistently improve in both lexical and syntactic naturalness.
@@
Our experiments do not include any low-resource languages because our approach relies on the availability of ground-truth distributions from native, human-written data, which is often scarce or unavailable for many languages. While our naturalness evaluation metrics are designed to be language-agnostic, they depend on reliable word tokenizers and dependency parsers for the languages studied. This is discussed in greater detail in Appendix G. Unfortunately, such tools are still lacking for most low-resource languages. However, we argue that naturalness evaluation should not be the priority for these languages at this stage, as it is only meaningful once models achieve a certain baseline level of overall performance.
@@
Acknowledgments
We would like to thank all the people at Apple who supported this work and provided helpful feedback. The majority of this work was carried out during Yanzhu Guo’s internship at Apple. Simone Conia gratefully acknowledges the support of the PNRR MUR project PE0000013-FAIR, which fully funds his fellowship since October 2023.
@@ Appendix C Implementation of Naturalness Metrics
Appendix D Comparison with N-gram Overlap Metrics
We compute BLEU and ROUGE scores for the models evaluated in the naturalness benchmark (Table 1) and report them in Table 7. These n-gram overlap metrics show markedly different behavior from our naturalness metrics. For example, while our metrics consistently identify the Qwen models as the least natural across all three languages, this pattern is not reflected in their BLEU or ROUGE scores. Similarly, our metrics capture a steady improvement in naturalness in newer LLM versions compared to earlier ones, a trend that BLEU and ROUGE fail to detect.
Appendix E Preference Dataset Generation
@@
E.1 Prompts
@@
E.2 Dataset Statistics
@@
Appendix F Hyperparameters for Preference Tuning
@@
Appendix G Adaptation of Naturalness Metrics for Other Languages
In our current approach, the lexical naturalness metric operates at the word level, as words are meaningful lexical units in the three languages (English, French and Chinese) analyzed in this study. However, for polysynthetic and agglutinative languages, which feature richer morphological structures and sparser word distributions, morphemes may be more suitable as the basic lexical units. However, adapting our metrics for these languages would require morphological parsing, which might be challenging in some cases, especially for polysynthetic languages. For the syntactic naturalness metric, we use POS units as defined by the UD grammar. In the languages we analyzed, these units typically align with individual words. However, in English and French, there exist multi-word tokens where a single orthographic token corresponds to multiple syntactic words, a detail which we will clarify in the final version of our paper. Extending our syntactic metrics to polysynthetic and agglutinative languages would similarly require using the specific annotation frameworks provided by the UD schema for each language. For instance, in Turkish, the UD schema also divides orthographic tokens into syntactic words. We thank the reviewer for raising this interesting question. We believe that future efforts to extend our framework to additional languages should address not only the resource disparity between high-resource and low-resource languages but also integrate linguistically informed decisions to account for the diversity of typological features across languages.